Solve the following differential equation

La solución de la ecuación diferencial y +4.2 Seleccione una: o cy=+C o d.y=PTER
Answer #1

y + 4. 4 = ex This is linear differential equation. of the form yt P(x) y = Q(x) Here, integrating factor (IF) If = eSP addx. S x y = sex xh da. now, we need to evaluate Cele de lett 85 dt da exs = ě (584).=) da= at exs (5x4) put nalue of t and da. dt

Your honest feedback is very important for better results.

Thanks